Magnesium hydroxide has a Ksp of 1.2 x 10^-11. What is the molar solubility of magnesium hydroxide, Mg(OH)2, when t is added to a 1.9 x 10^-2 M solution of NaOH?
Answer #1

Mg(OH)2(s) ----------------> Mg^2+ (aq) + 2OH^- (aq)

                                            s                  2s+1.9*10^-2

Ksp     = [Mg^2+][OH^-]^2

1.2*10^-11   = s*(2s+1.9*10^-2)                             [ 2s + 1.9*10^-2   = 1.9*10^-2]

1.2*10^-11     = s*1.9*10^-2

s                    = 1.2*10^-11/1.9*10^-2    = 6.3*10^-10M

solubility of Mg(OH)2   = 6.3*10^-10M